OpenStudy (anonymous):

solve by substitution -6x=-9-y 2y-12x=18

OpenStudy (jdoe0001):

if you were to solve for "y" -6x=-9-y what would that look like?

OpenStudy (anonymous):

can you explain the steps

OpenStudy (jdoe0001):

-6x=-9-y # add +y to each side -6x + y = -9-y+y => -6x+y = -9 # add 6x to each side -6x+y+6x = -9+6x => ?

OpenStudy (anonymous):

(2,3)

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

well, is not that, what would the equation give you at # add 6x to each side -6x+y+6x = -9+6x => ????

OpenStudy (anonymous):

(3,9) ?

OpenStudy (jdoe0001):

you'd firstly need to solve for "y" -6x+y+6x = -9+6x

OpenStudy (jdoe0001):

so, cancelling like terms there would give you, a "y" you can use

OpenStudy (anonymous):

no solution? thats my last option

OpenStudy (jdoe0001):

or your can solve for "y" on that one above